Loan amortization and EAR

Loan amortization and EAR

You want to buy a car, and a local bank will lend you $25,000. The loan will be fully amortized over 5 years (60 months), and the nominal interest rate will be 6% with interest paid monthly.

  1. What will be the monthly loan payment? Round your answer to the nearest cent.

    $  

  2. What will be the loan's EAR? Round your answer to two decimal places.

     %
Answer #1

Using the financial calculator, we set the below:

PV=25000

N=60

I/Y=6%/12=0.5%

CPT PMT=$483.32

Loan's EAR=(1+6%/12)^12-1=6.1677%


answered by: Invious